40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jun 16 23:15:44 2014, Jenne, Update, LSC, IFO alignment recovery Note_061614_230723_0.jpg
    Reply  Tue Jun 17 17:14:14 2014, Manasa, Update, LSC, IFO alignment recovery 
       Reply  Tue Jun 17 21:49:13 2014, Jenne, Update, LSC, IFO alignment recovery 
    Reply  Wed Jun 18 11:57:44 2014, not Jenne, Update, LSC, IFO alignment recovery 
Message ID: 10055     Entry time: Wed Jun 18 11:57:44 2014     In reply to: 10047
Author: not Jenne 
Type: Update 
Category: LSC 
Subject: IFO alignment recovery 

Quote:

I noticed today, and Rana said that he saw Saturday, that the MC refl value when the MC is unlocked is unusually high.  It typically goes to about 4.5 V, but now is going up to 6.5V.  Since the PMC output is the same as usual (max seen has been about 0.82 today), something must have happened between the PMC and the IMC. 

Late last week, EricG and Nichin were looking at things on the AS table.  Was anything touched on the AS table?  Was anything touched on the PSL table?  'Fess up please, so that we can pinpoint what the change was.

 

 Nope, we did not touch any of the PDs other than AS55. I have mentioned in  my elog:10037 what we did exactly.

We just looked at all the other PDs to check if they were being illuminated by the correctly labeled fiber. Nothing other than that.

ELOG V3.1.3-